solve for the given variables
4/(x+2)= 16/(x+5)

Answer by tommyt3rd(5050)

4/(x+2)= 16/(x+5)
1/(x+2)=4/(x+5)
x+5=4(x+2)
x+5=4x+8
-3=3x
x=-1